OpenStudy (anonymous):

The diameter of a basketball rim is 18 inches. A standard basketball has a circumference of 30 inches. What is the distance between the ball and the rim in a shot in which the ball goes exactly in the center of the rim? Show your work. My stuff so far: The required value = ( 18 - d)/2 where d = diameter of the basketball. What else do I put down??

OpenStudy (anonymous):

total room for the ball=18-9.55=8.45inches if the ball goes through exactly in the center then there will be 8.45/2=4.225inches of room between rim and ball. Is that the correct answer?

OpenStudy (anonymous):

(18 - 9.55)/2 = 4.2in, inches between the ball and the rim on each side? Is that it?

OpenStudy (phi):

looks good. the answer depends on how many digits they want

OpenStudy (anonymous):

Is it in order?? Is it neat enough? And do I have the answer?

OpenStudy (anonymous):

I need to write it down, step by step, in order right down to the answer

OpenStudy (phi):

It should have a figure with colors and lines and labels and footnotes. kidding... it looks good.

OpenStudy (mathstudent55):

I would only add the actual calculation of the diameter of the ball based on its diameter.

OpenStudy (anonymous):

circumference (C) = 2(pi)r \( \huge \frac{C}{2\pi} = r \) \( \huge \frac{30}{2\pi} = r \) \( \huge r = \frac{30}{2\pi}\) \( \huge r = 4.77\) \( \huge d = 4.77 \times 2\) \( \huge d = 9.54\)

OpenStudy (anonymous):

18 - 9.54 = 8.46

OpenStudy (anonymous):

Thank you guys xD

OpenStudy (anonymous):

oh btw. @phi @mathstudent55 is nixy correct?

OpenStudy (anonymous):

8.46 / 2 = 4.23

OpenStudy (anonymous):

@Lollygirl217 I am correct.

OpenStudy (mathstudent55):

Yes

OpenStudy (mathstudent55):

But I prefer the approximation d = 9.55 in.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Thank you nixy

OpenStudy (anonymous):

YW